# A General Bayesian Framework for Informative Input Design in System Identification

### v1

@@ IV-D Algorithm
Notice that when computing $x_{n+1}$ in eq. 11 we choose $\Sigma_{i}\leftarrow\hat{\Sigma}_{+}$ from (13) for $i=1,\dots,n+1$. When performing the maximum a posteriori estimation in eq. 8 , however, we use $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$. This selection is made because our next input should excite the system in a way that is informative and decreases the uncertainty present in our current system estimate. However, when improving the estimate of the model parameters we should use less informative covariances to robustify the estimation.
Algorithm 1 Proposed Algorithm for Parameter Estimation and Input Design
1:dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, parameter estimate $\hat{\theta}$, covariances $\Sigma_{\mathrm{prior}}$ and $\Sigma_{\mathrm{meas,prior}}$, model family $f$, constraint sets $\mathcal{X},\mathcal{Y}$, constants $\delta,\lambda$ and $w$
2:$\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}},\ \mathrm{for\ all}\ i=1,\dots,n$
@@
4:$\Sigma_{i}\leftarrow\hat{\Sigma}_{+},\ \mathrm{for\ all}\ i=1,\dots,n+1$
5:Obtain next input $x_{n+1}$ by solving eq. 11
6:Obtain output $y_{n+1}$ for input $x_{n+1}$ using the true system
7:return dataset $\{(x_{i},y_{i})\}_{i=1}^{n+1}$
@@ V Results
We test our algorithm on both linear and nonlinear systems. [^1] The code is included in the repository: https://github.com/alextzik/informative_input_design We further perform an ablation study to show the importance of finetuning the covariances $\Sigma_{i}$ in our algorithm.
@@
Our emphasis is on the informative input design. We compare against three different model-free baselines for the selection of inputs. We maintain the same parameter estimation module from section IV in all methods. The first baseline is random. At every iteration, we run algorithm 1, with a change in line 4: we suppose $x_{n+1}\sim\mathcal{N}(0,I)$, i.e., $x_{n+1}$ is randomly selected. The second baseline is based on pseudorandom binary sequences and is called PRBS. At every iteration, we run algorithm 1, with a change in line 4: we suppose that each coordinate of $x_{n+1}$ originates from a pseudorandom binary sequence. Finally, the last baseline is called mutlisine and is based on orthogonal multisine inputs [10]. In line 4 of algorithm 1, we set the coordinate $i$ of $x_{n+1}$ as follows
$x_{n+1,i}=\sum_{j=0}^{M-1}A_{i,j}\sin\left(\dfrac{2\pi(d_{x}j+i)(n+1)}{T}+\phi_{i,j}\right),$
where $M$ is the total number of sinusoidal components and $T$ is the total number of inputs to be designed. We sample each phase $\phi_{i,j}$ uniformly in $\left[0,2\pi\right]$ and each amplitude $A_{i,j}$ uniformly in $\left[-1,1\right]$. We select $M=30$ in all experiments below.
Algorithm Parameters
We set $w=0.2$, $\delta=0.3$, and $M=30$. We do not consider any constraint s on $\mathcal{X}$ or $\mathcal{Y}$ for simplicity in the exposition. We do set an upper bound on the allowed input magnitude. If the input provided by any method has a larger magnitude, it is scaled to have the largest allowed magnitude. The first datapoint is assumed fixed for all methods.
@@
In all cases below , the true system belongs to the parametric model family. We can thus compare the methods with respect to the error $\lVert\hat{\theta}-\theta_{\mathrm{true}}\rVert_{\infty}$ as a function of iteration number. Further, we can compare the methods with respect to the $\log\det$ of the $\Sigma_{\mathrm{posterior}}$ at the final timestep, which is given in eq. 9. We assume $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$. A lower value signifies less uncertainty in the parameter estimate and a more informative sequence of inputs. The results are gathered in table I for all experiments.
@@ V-A Linear System
Our results are shown in Figure 2. In addition to the error, we plot the trajectory of the parameter estimate along with its error ellipses. We only show the estimate d trajectory of the first two coordinates of $\hat{\theta}$, although the complete estimate is 4-dimensional. We further plot the designed inputs for the various methods. Our algorithm exhibits a faster convergence to the true model parameters compared to the baselines. This is verified by the lower value of $\log\det$ for the final error covariance.
@@ V-B Hénon map dynamics
Our results are shown in Figure 3. Our proposed algorithm is able to converge to the true model parameters and outperforms the baselines with respect to both metrics considered. From the estimate trajectory, we observe that the covariance truly captures the uncertainty in the parameter: based on the true parameter the estimate has a higher uncertainty in the second coordinate $\theta_{2}$, which is captured by the elongated ellipse axes along that direction.
@@ V-C Unicycle Model
The time discretization is thus one of the unknown parameters. We start the model from the origin with zero being the first control for all methods. Figure 4 shows the performance of our algorithm. It outperforms the considered baselines. We further plot $(x_{t,1},x_{t,2})$ for the various methods.
V-D Ablation
In order to show the significance of finetuning, we compare, in the case of the Hénon dynamics, algorithm 1 with a variant that uses $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$ throughout. The new variant is termed proposed (no finetuning). In this simulation , we remove the amplitude constraint. It is evident in Figure 1 that the variant does not converge properly. This highlights the importance of finetuning the covariances $\Sigma_{i}$.
Figure 1: Ablation study in the case of Hénon dynamics.
(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
(b) Model parameter estimate error over timesteps.
(c) Inputs as a function of timestep for the various methods.
@@ V-D Ablation
(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
(b) Model parameter estimate error over timesteps.
(c) Inputs as a function of timestep for the various methods.
@@
(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
(b) Model parameter estimate error over timesteps.
(c) State of dynamics as a function of timestep for the various methods.
@@
proposed $\mathbf{-18.5}$ $\mathbf{-14.9}$ $\mathbf{-9.8}$
random $-10.7$ $-5.9$ $-7.0$
PRBS $-10.2$ $-5.1$ $-9.2$
multisine $-13.7$ $-9.7$ $-8.9$
@@ VI Conclusion
We present a Bayesian methodology for informative input design that is based on linear Gaussian assumptions. It extends to general nonlinear system identification through the Taylor expansion. A key module in our algorithm is the online calibration of the noise covariances that greatly impacts performance. Our algorithm also has limitations. We must select the trust region shape and volume when performing the parameter estimation , and we do not address the problem of selecting the parametric model family. Future work will propose a method for judging the suitability of a given parametric model family.

### v2

@@ IV-D Algorithm
Our algorithm first runs a loop to improve the parameter estimate. We obtain a new parameter estimate by solving problem (8). We only accept the new solution if the linearization errors are insignificant compared to the model errors. In this case, we update the covariance $\Sigma$ and resolve problem (8) using the new covariance and the Taylor expansion around the new parameter estimate. If the linearization errors are deemed significant we reject the new estimate, reduce $\delta$ and resolve. This iterative technique will produce a new parameter estimate. It will further produce an accurate error covariance $\Sigma_{\mathrm{model\ error}}$ for the errors attributed to the discrepancy between the current model and the true system.
In the second phase, we obtain the next input by optimizing eq. 11. Note that we use $\Sigma_{\mathrm{model\ error}}$, because we aim to reduce the uncertainty associated with the model mismatch.
The algorithm outputs an updated $\Sigma$, which is used as input in the next algorithm call.
Algorithm 1 allows for testing whether the chosen parametric model family $f$ is suitable for the system under consideration. We know that the parameter estimates obtained from the successive calls to algorithm 1 correspond to the best model fits for the obtained datapoints and Taylor MAP objective. Suppose that the magnitude of $\Sigma_{\mathrm{model\ error}}$ does not decrease between successive calls to algorithm 1. This suggests that the chosen parametric family $f$ is not appropriate for the system.
Algorithm 1 Proposed Algorithm for System Identification
1:dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, parameter estimate $\hat{\theta}$, covariances $\Sigma_{\mathrm{prior}}$ and $\Sigma$, model family $f$, constraint sets $\mathcal{X},\mathcal{Y}$, constants $\delta$ and $\lambda$
2:repeat
@@
4:     Compute $\hat{\epsilon}_{i},\bar{\epsilon}_{i}$ for all $i=1,\dots,n$
5:     if $n^{-1}\sum_{i=1}^{n}\lVert\bar{\epsilon}_{i}\rVert$ large compared to $n^{-1}\sum_{i=1}^{n}\lVert\hat{\epsilon}_{i}\rVert$ then
6:         Reduce $\delta$
7:     else
8:         Update $\Sigma$ using eq. 16
9:         Set $\Sigma_{\mathrm{model\ error}}=\dfrac{1}{n}\sum_{i=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top}$
10:         $\hat{\theta}\leftarrow\hat{\theta}_{+}$
11:until convergence
12:Obtain next input $x_{n+1}$ by solving eq. 11 where $\Sigma\leftarrow\Sigma_{\mathrm{model\ error}}$
13:Obtain output $y_{n+1}$ for input $x_{n+1}$ using the true system
14:return dataset $\{(x_{i},y_{i})\}_{i=1}^{n+1}$, $\hat{\theta}_{+}$, $\Sigma$, and $\Sigma_{\mathrm{model\ error}}$
@@ V Results
We test our algorithm on both linear and nonlinear systems. [^1] The code is included in the repository: https://github.com/alextzik/informative_input_design Our results show the importance of finetuning the covariances $\Sigma$ in our algorithm.
@@
We consider a variant of our proposed algorithm as baseline. The baseline does not rely on any finetuning. Given the dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, $\Sigma$, $\Sigma_{\mathrm{prior}}$ and $\hat{\theta}$, it obtains an updated estimate $\hat{\theta}_{+}$ by solving eq. 8. It then solves eq. 11 to obtain the next input $x_{n+1}$. It does not update $\Sigma$, which is always equal to the initial value provided.
Set-Up
We set $\delta=0.3$. In line 5 of algorithm 1, we reduce $\delta$ by setting $\delta\leftarrow 0.8\delta$. We do not consider any constraint s on $\mathcal{X}$ or $\mathcal{Y}$ for simplicity in the exposition. The first datapoints are assumed fixed for all methods. Further, we run the algorithms with multiple initializations of $\Sigma$. We use the scalar function $\log\det$ as $\mathcal{M}$ when determining the next input.
@@
In most cases below , the true system belongs to the parametric model family. We can thus compare the methods with respect to the error $\lVert\hat{\theta}-\theta_{\mathrm{true}}\rVert_{\infty}$ as a function of iteration number. Further, we can compare the methods with respect to the $\log\det$ of the $\Sigma_{\mathrm{posterior}}$ at the final timestep, which is given in eq. 9. A lower value signifies less uncertainty in the parameter estimate and a more informative sequence of inputs. The results are gathered in Table I for all experiments.
@@ V-A Linear System
Our results are shown in Figure 3. In this experiment, we also bound the norm of the input to be less than $0.5$. In addition to the error, we plot the trajectory of the parameter estimate along with its error ellipses. We only show the estimate d trajectory of the first two coordinates of $\hat{\theta}$, although the complete estimate is 4-dimensional. We further plot the designed inputs for the two methods. Our algorithm exhibits a faster convergence to the true model parameters compared to the baseline. This is verified by the lower value of $\log\det$ for the final error covariance.
@@ V-B Hénon map dynamics
Our results are shown in Figure 4. Our proposed algorithm is able to converge to the true model parameters and outperforms the baseline with respect to both metrics considered. From the estimate trajectory, we observe that the covariance truly captures the uncertainty in the parameter: based on the true parameter value, the estimate along its trajectory has a higher uncertainty in the first coordinate $\theta_{1}$, which is captured by the elongated ellipse axis along that direction. We also observe that the error ellipses shrink as we move towards the true parameters. This implies that our uncertainty is reduced.
@@ V-C Unicycle Model
The time discretization is thus one of the unknown parameters. We start the system from the origin with zero being the first control for both methods. Figure 5 shows the performance of our algorithm. It outperforms the considered baseline. The posterior error ellipses are an accurate representation of the uncertainty. Once, $\theta_{1}$ is correctly estimated, the ellipses become vertical, since there is uncertainty only in $\theta_{2}$. We further plot the state trajectory $(x_{t,1},x_{t,2})$ for the two methods. In this experiment, we bound the norm of the input to be less than $1$.
V-D Importance of Finetuning
In order to show the significance of finetuning, we compare, in the case of the Hénon dynamics, algorithm 1 with the baseline. The initial covariance $\Sigma$ is set to $0.2I$, while $\Sigma_{\mathrm{prior}}=0.1I$ and $\theta_{\mathrm{prior}}=(7,4)$. It is evident in Figure 1 that the baseline does not converge properly. This can be attributed to the high confidence we have initially placed on both the prior value of $\theta$ and the accuracy of the measurements. Under these conditions, without finetuning, it is impossible for the algorithm to detect the false confidence. This highlights the importance of finetuning the covariance $\Sigma$.
Figure 1: Importance of finetuning the covariance matrix $\Sigma$ in the case of Hénon dynamics.
V-E Model Mismatch
We consider the case where the parametric model family does not include the true system. We suppose that the true system is the Hénon dynamics of eq. 19 and the parametric model family consists of linear models of the form
$f(x_{t};\theta)=\begin{bmatrix}\theta_{1}&\theta_{2}\\ \theta_{1}&\theta_{2}\end{bmatrix}x_{t}.$ (23)
Obviously, this is an inadequate model family. As discussed in section IV-D, we expect our algorithm to provide an indication of the inadequacy of the parametric model family via the magnitude of $\Sigma_{\mathrm{model\ error}}$. This is shown in Figure 2. The magnitude of $\Sigma_{\mathrm{model\ error}}$, here chosen as $\log\det\Sigma_{\mathrm{model\ error}}$ does not decrease as the iterations progress. This implies that the realized errors remain high. Given that our algorithm computes MLEs, this indicates that the obtained parameter estimates, which are in a sense optimal, are inadequate to capture the true dynamics.
Figure 2: Mean $\log\det\Sigma_{\mathrm{model\ error}}$ and 1-standard deviation interval over iterations when the parametric model family cannot capture the true system dynamics. The fact that $\log\det\Sigma_{\mathrm{model\ error}}$ does not decrease indicates an inadequate parametric model family.
(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
(b) Inputs as a function of time for the various methods for a specific initialization of $\Sigma$.
(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
@@ V-E Model Mismatch
(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
(b) Inputs as a function of time for the various methods for a specific initialization of $\Sigma$.
(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
@@
(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
(b) State of dynamics as a function of time for the various methods for a specific initialization of $\Sigma$.
(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
@@
proposed $\mathbf{-77.2}$ $\mathbf{-51.2}$ $\mathbf{-43.5}$
baseline $-12.1$ $-31.8$ $-7.5$
@@ VI Conclusion
We present a Bayesian methodology for system identification that is based on linear Gaussian assumptions. It extends to general nonlinear system identification through the Taylor expansion. A key module in our algorithm is the online calibration of the noise covariance that greatly impacts performance and allows us to judge the suitability of a parametric model family for the underlying system. Our algorithm also has limitations. The use of the trust region, although necessary to ensure the accuracy of the Taylor series, might forbid large changes in the parameters when these are implied by the data. Future work will apply our method to high-dimensional problems and look into theoretical guarantees.